### v1

@@ 6 Data quality monitoring
To ensure adequate data quality during the deployment and commissioning phases of GP300 and G@A, a few key parameters are tracked in every DU. As previously stated, a complete event consists of the triggered ADC time traces and the slow control parameters, like the battery level at the board input, the GPS, and, in G@A alone, the ADC and FPGA chip temperatures and the trigger rate.
@@
Through the unbiased triggered data, these parameters can be continuously monitored. Their behavior was used not only to commission the arrays but also to validate the firmware and the local and central DAQ software. Other features—such as their root mean square (RMS), fast Fourier transform, and power spectral densities (PSDs)—are extracted from the ADC traces in the time domain to monitor day-night effects and detect local radio transients. Particularly, the PSDs can be used to characterize the environment radio background, the electronic noise self-induced by the setup, the Galactic radio background, and the RFI. These parameters are checked daily to ensure optimal operation of the setups. We monitor their behavior over time, comparing them between DUs to achieve consistent data quality across the array.
@@
The commissioning is facilitated by a data monitoring web page. The page is updated daily at 05:00 CET, by a crontab job that runs a series of scripts at CC-IN2P3 on the data from both arrays. The DU uptime, the mean RMS of the last 24 hours of traces, battery levels, and temperatures are tracked as a function of time. Averaged PSDs are also displayed for easy comparison between channels and between DUs. The page also stores the averages over the last 7 days and the last 30 days. There is flexibility to present the data of all DUs together or of only the desired ones.
7 First results
@@
Figure 15 shows spectrograms generated using data collected by two DUs, one from GP300 and one from G@A, over a period of about nine and thirteen days, respectively. These spectrograms are representative of the standard behavior of all DUs in their respective arrays.
For GP300, the PSDs are largely free from strong narrowband noise in all polarizations, except within 118–140 MHz, where there is identified man-made noise (cf. Fig. 16). Figure 15, left, shows an apparent daily modulation in the PSD, although further studies are required to properly identify it as a day-night effect or as the diffuse Galactic radio emission.
For G@A, several narrowband noise sources are present in the three polarizations, some continuous throughout the data acquisition time and some not. Of the intermittent contributions, one at around 70 MHz can be seen in all polarizations, where there is an increase of amplitude for a few hours with no clear periodicity. In the North-South polarization, the four lines within 60–80 MHz correspond to the known emission of the AERA beacon, set up at the CRS (see Fig. 2).
In Fig. 15, the white horizontal lines represent periods of DU hibernation, after which data acquisition resumes autonomously once the conditions are appropriate. The G@A DU had more hibernation periods than the GP300 DU, a reflection of the harsher environment in Argentina, leading to more periods with insufficient sunlight to power the batteries and more instances of overheating. In general, the radio environment at GP300 is considerably quieter than at G@A, but wide background-free frequency bands exist at both sites. Nevertheless, the background in several frequency ranges is substantial and can negatively influence the self-trigger capabilities of the system, especially in G@A.
@@
Figure 16 shows the time-averaged PSDs of GP300 and G@A, where the background noise contributions can be more clearly seen. For GP300, the PSD is mostly free of peaks, except in the 118–140 MHz range—in agreement with Fig. 15—where there are strong contributions from plane and satellite communications. These sources have high intensity and spread over multiple frequency bins, collectively defining a frequency band that could be excluded in analyses. Above 140 MHz, there is a slight drop in the PSD baseline level. Since this range is dominated by internal noise, only gain effects are expected, which is consistent with the drop seen in the system performance shown in the transfer functions in Fig. 12.
@@ 83.2 NS TV channel 6, video
For G@A, we carried out a dedicated run with longer time series and finer frequency resolution (see the caption of Fig. 16) to resolve individual noise sources more clearly. In Fig. 16, the drop-off in the PSD at low frequencies is due to the impedance mismatch between the LNA and the antenna.
Tables 1 and 2 show the identified sources of several of the peaks of the GP300 and G@A spectra in Fig. 16. In particular, the G@A PSD characterizes the electromagnetic background at the Pierre Auger site at frequencies much higher than either AERA or the Auger Radio Detector are capable of measuring. This can aid and complement analyses of the Pierre Auger Collaboration—another example of the synergy between the Pierre Auger Observatory and GRAND.
Figures 15 and 16 show that, while radio background is unavoidable in both arrays, it is localized within frequency ranges that can be isolated and masked out in specific analyses. For instance, searches for the diffuse Galactic radio emission require noise-free frequency ranges, due to the lower emission amplitude of this background compared to artificial radio sources. This search is ongoing in both prototype arrays [29, 30], since it presents an opportunity for detector calibration.
The above results show that the radio background at both sites is understood. At the GP300 site, there are fewer RFI sources compared to the G@A site. The baseline levels of the PSDs are the same for both prototype arrays, around $10^{-2}$ mV2 MHz-1. In summary, early data from the GRAND prototype arrays already validate their usefulness: that, despite differences in hardware and software, the overall response of the GRAND DU is consistent regardless of the deployment site. This is a structural finding for future cross-calibration runs and for the later stages of the experiment, which envision two or more 10,000-DU arrays deployed at different geographical locations, and whose data will be analyzed jointly.
@@ 8 Summary and perspectives
The Giant Radio Array for Neutrino Detection (GRAND) is a planned observatory of ultra-high-energy particles of cosmic origin: cosmic rays, neutrinos, and gamma rays. GRAND aims to detect the radio emission from air showers initiated by their interaction in the atmosphere using large ground arrays of antennas. The next large deployment goal of GRAND is to have two arrays consisting of 10,000 antennas each in the Northern and Southern Hemispheres. To achieve this, the GRAND Collaboration has adopted a staged approach to deployment, with three ongoing prototype arrays to test and validate the detection and reconstruction methods.
@@
Three prototype arrays were deployed in 2023. GRAND@Nançay, in France, is a 4-antenna array used as a testbench for the Collaboration’s European laboratories. GRANDProto300, in China, has gone through two stages of deployment, from an initial 13-antenna stage to its current 65-antenna stage. These stages act as pathfinders for the future final 300-antenna array. GRAND@Auger, in Argentina, stems from an agreement between the GRAND and the Pierre Auger Collaborations. The 10-antenna array was deployed on modified AERA stations. In this paper, we focused on the hardware, software, and early operation of GRANDProto300 and GRAND@Auger.
To adapt to local conditions and explore optimal designs for the final GRAND setup, the prototype array at each site featured different key choices in hardware and software. Throughout 2023, the arrays, including their hardware, firmware, and DAQ software, were commissioned using daily data monitoring. Operating these prototypes under prolonged field conditions exposed hardware and software challenges that have been met.
GRANDProto300 and GRAND@Auger have remained largely operational for about two years. Their data have been analyzed and understood. We have shown results based on these data that demonstrate the adequacy of the GRAND experimental setup and provide an understanding of the radio background present at each site. The detection of the Galactic radio emission—the search for which is ongoing in both prototypes—will refine the calibration of the GRAND detection units.
The GRAND prototypes have already yielded a key result: that the performance of the GRAND detection units is similar at the two array sites, despite one having a radio background notably higher than the other, and despite their differences in hardware and software. The demonstrated resilience of the GRAND DU is not merely a technical achievement; it is a foundational scientific result for the entire project. A future stage of GRAND consisting of arrays distributed in multiple geographical locations will inevitably face diverse environmental and radio-background conditions. Our finding that consistent performance can be achieved through adaptable hardware and software confirms the feasibility of combining data from geographically distant arrays. This is essential for achieving the full-sky coverage and large-scale exposure of future stages of GRAND.
The ultimate goal of GRANDProto300 and GRAND@Auger is to validate the autonomous radio-detection of inclined extensive air showers in GRAND, expected from ultra-high-energy cosmic rays, neutrinos, and gamma rays. This detection, triggered by no signal other than radio, and with a high efficiency, is essential to the goals of GRAND. Presently, the data collected by GRANDProto300 and GRAND@Auger are being searched for air showers initiated by ultra-high-energy cosmic rays. In GRANDProto300, we have identified the first cosmic-ray candidates [38, 39, 40, 41, 42, 43]. In GRAND@Auger, the detection of cosmic-ray candidates is being validated through their coincident detection by the Pierre Auger Observatory [47].
@@ Acknowledgments
The GRAND Collaboration acknowledges the significant contributions to this manuscript of Beatriz de Errico, for her work on data management and analysis, site survey and construction, DAQ software debugging, manuscript writing, and figure preparation; and of Shen Wang, for his work on GP300 site survey and construction, hardware verification, and manuscript planning and writing. In addition, the Collaboration acknowledges significant contributions from the following Collaboration members: Mauricio Bustamante (coordination, manuscript writing, figures), Nicoleta Cucu Laurenciu (firmware development and trigger algorithm implementation), Bohao Duan (construction, GP300 DAQ software development), François Legrand (section on data storage), Pengxiong Ma (construction, RFI identification, GP300 data management and analysis), Frédéric Magnard (communications at GP300, DAQ and communications at Auger), Stavros Nonis (data analysis, RF chain implementation, and figures), Lech Piotrowski (GP300 DAQ software), Daniel Szálas Motesiczky (LNA design, electronics validation), Xishui Tian (laboratory trigger tests), Anne Timmermans (data monitoring web page), Charles Timmermans (site survey, construction, DAQ software development and debugging, data management and analysis, manuscript writing and revising, figures), Feng Wei (LNA design), Xin Xu (antenna response, RF chain optimization, data analysis), Xing Xu (construction, GP300 hardware verification, firmware development and trigger algorithm implementation), Pengfei Zhang (site survey, redesign and functional test of detection unit, construction, data analysis, RF chain generation), Yi Zhang (site survey and construction, data analysis, manuscript writing). The GRAND Collaboration is grateful to Arno Engels (Radboud University) for the mechanical design of the detection units, to Fei Gao (Xidian University) for that and for the thermal design of the detection units, to the local government of Dunhuang during site survey and deployment approval, to Yu Tang for his help on-site at the GRANDProto300 site, and to the Pierre Auger Collaboration, including the staff in Malargüe, for the warm welcome and continuing support, and the Auger Publications Committee, especially Silvia Mollerach, for their feedback. The GRAND Collaboration acknowledges the support from the following funding agencies and grants. Brazil: Conselho Nacional de Desenvolvimento Cientiífico e Tecnológico (CNPq); Fundação de Amparo à Pesquisa do Estado de Rio de Janeiro (FAPERJ); Coordenação Aperfeiçoamento de Pessoal de Nível Superior (CAPES). China: NAOC, National SKA Program of China (Grant No. 2020SKA0110200); Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061); Program for Innovative Talents and Entrepreneurs in Jiangsu, and High-end Foreign Expert Introduction Program in China (No. G2023061006L); China Scholarship Council (No. 202306010363). Denmark: Villum Fonden (project no. 29388). France: “Emergences” Programme of Sorbonne Université; France-China Particle Physics Laboratory; Programme National des Hautes Energies of INSU; for IAP—Agence Nationale de la Recherche (“APACHE” ANR-16-CE31-0001, ANR-23-CPJ1-0103-01), CNRS Programme IEA Argentine (“ASTRONU”, 303475), CNRS Programme Blanc MITI (“GRAND” 2023.1 268448), CNRS Programme AMORCE (“GRAND” 258540); Fulbright-France Programme; IAP+LPNHE—Programme National des Hautes Energies of CNRS/INSU with INP and IN2P3, co-funded by CEA and CNES; IAP+LPNHE+KIT—NuTRIG project, Agence Nationale de la Recherche (ANR-21-CE31-0025); IAP+VUB: PHC TOURNESOL programme 48705Z. Germany: NuTRIG project, Deutsche Forschungsgemeinschaft (DFG, Projektnummer 490843803); Helmholtz—OCPC Postdoc-Program. Poland: Polish National Agency for Academic Exchange within Polish Returns Program no. PPN/PPO/2020/1/00024/U/00001,174; National Science Centre Poland for NCN OPUS grant no. 2022/45/B/ST2/0288. USA: U.S. National Science Foundation under Grant No. 2418730. Computer simulations were performed using computing resources at the CC-IN2P3 Computing Centre (Lyon/Villeurbanne, France), partnership between CNRS/IN2P3 and CEA/DSM/Irfu, and computing resources supported by the Chinese Academy of Sciences.

### v2

@@ 5.2 Quality monitoring
To ensure adequate data quality during the deployment and commissioning phases of GP300 and G@A, a few key parameters are tracked in every DU. A complete event consists of the triggered ADC time-traces and the slow control parameters, such as the battery level at the board input, the GPS information, and, in G@A alone, the ADC and FPGA chip temperatures and the trigger rate.
@@
Through the unbiased triggered data, these parameters can be continuously monitored. Their behavior was used not only to commission the prototype arrays but also to validate the firmware and the local and central DAQ software. Other features—such as their Root Mean Square (RMS), Fast Fourier Transforms (FFTs), and Power Spectral Densities (PSDs)—are extracted from the ADC traces in the time domain to monitor day-night effects and detect local RFI transients. In particular, the PSDs can be used to characterize the environment radio-frequency background, the thermal noise from the electronics, and the Galactic noise. These parameters are checked daily to ensure optimal operation of the prototypes. We monitor their behavior over time, comparing them between DUs to achieve consistent data quality across the array.
@@
The commissioning is facilitated by a data monitoring web page. The page is updated daily at 05:00 CET, by a crontab job that runs a series of scripts at CC-IN2P3 on the data from both arrays. The DU uptime, the mean RMS of the last 24 hours of traces, battery levels, and temperatures are tracked as a function of time. The averaged PSDs are also displayed for easy comparison between channels and between DUs. The page also stores the averages over the last 7 days and the last 30 days. There is flexibility to present the data of all DUs together or of only the desired ones.
6 Background and noise characterization
@@
Figure 14 shows spectrograms generated using unbiased triggered data collected by two DUs, one from GP300 and one from G@A, over a period of about nine and twelve days, respectively. These spectrograms were calculated from voltage time-series measured at the ADC input and are representative of the standard behavior of all DUs in their respective arrays. For GP300, the PSDs are largely free from strong narrowband RFIs in all polarizations, except within 118–140 MHz, where anthropogenic background is identified. In addition, Figure 14, left, shows an apparent daily modulation in the PSD.
For G@A, several broad and narrowband RFIs are present in the three polarizations, some continuous throughout the data acquisition time and some not. Of the intermittent contributions, one at around 70 MHz can be seen in all polarizations, where there is an increase of amplitude for a few hours with no clear periodicity. In the North-South polarization, the four lines within 60–80 MHz correspond to the known emission of the AERA beacon, set up at the CRS (see Fig. 2). The higher variability of background and noise at G@A compared to GP300 is reflected in the larger granularity of the spectrograms.
In principle, unbiased data are collected with a 100% duty cycle, unless the acquisition is interrupted. For instance, in Fig. 14, the white horizontal lines correspond to periods of DU hibernation, in which data acquisition was suspended and then resumed autonomously once conditions are appropriate. The G@A DU had more hibernation periods than the GP300 DU, a reflection of the harsher environment in Argentina, leading to more periods with insufficient sunlight to power the batteries and more instances of overheating. In general, the electromagnetic environment at GP300 is considerably quieter than at G@A, but there are RFI-free frequency bands at both sites. Nevertheless, the background in several frequency ranges is substantial and can negatively influence the self-trigger capabilities of the system, especially in G@A.
@@
Figure 15 shows the time-averaged PSDs of GP300 and G@A, where the background RFI contributions can be seen more clearly. Similarly to the spectrograms, unbiasedly triggered voltage traces, also measured at the ADC output, were used to evaluate the average PSDs.
@@ 83.2 NS TV channel 6, video
For GP300, the PSD is mostly free of peaks, except in the 118–140 MHz range where there are strong contributions from plane and satellite communications. These sources have high intensity and are spread over multiple frequency bins, collectively defining a frequency band that will be excluded in analyses. Above 140 MHz, there is a slight drop in the PSD baseline level. Since this range is dominated by electronic noise, only gain effects are expected, which is consistent with the drop seen in the system performance shown in the transfer functions in Fig. 10. For G@A, we performed a dedicated run with longer time-series and finer frequency resolution (see the caption of Fig. 15) to resolve individual background RFI sources more clearly. In Fig. 15, the drop in the PSD at low frequencies is due to the impedance mismatch between the LNA and the antenna.
Tables 1 and 2 show the identified sources of several of the peaks of the GP300 and G@A spectra in Fig. 15. In particular, the G@A PSD characterizes the electromagnetic background at the Pierre Auger site at frequencies much higher than either AERA or the Pierre Auger Radio Detector are capable of measuring. This can aid and complement analyses of the Pierre Auger Collaboration—an example of the synergy between the Pierre Auger Observatory and GRAND.
Figure 16: Galactic synchrotron emission measurement with GRAND detection units. Power Spectrum Densities (PSDs), built from the East-West polarization, were averaged over a period of two months. Left: For GP300, unbiased data from June and July 2025 collected by DU 6 were used, in the frequency range of 60–80 MHz. Right: For G@A, unbiased data from DU 83 were also used in the frequency range of 100–200 MHz. The acquisition period was from March to June 2024, which, after a quality selection, summed up to 50 days. A amplitude maximum around the expected peak Galactic emission is clearly visible for both prototypes.
Identifying the frequency of narrowband RFIs facilitates their exclusion from the measured time-traces. In this way, we can obtain large RFI-free frequency bands, despite the high background content at both sites. This allows us to perform a first characterization of the stochastic noise level at each prototype array. The two main contributions expected within the GRAND working frequency range are thermal noise from the electronics and the Galactic synchrotron emission. The latter originates mainly from the Galactic plane, yet it is brightest in the inner Galaxy region. Thus, the emission reaches peak amplitude when the detector field of view is turned towards the Galactic center. If the GRAND DU thermal noise does not dominate the measurements, we expect to observe an increase in the average PSD between 16 and 20 hours in Local Sidereal Time (LST) corresponding to the Galactic emission.
Figure 16 shows the average PSDs for limited frequency ranges—60–80 MHz for GP300 (left) and 100–200 MHz for G@A (right)—. For both setups, the data was averaged over a period of about two months, sufficient for a significant observation of the Galactic noise. The differences in frequency range are a reflection of the different optimization frequencies of the LNA design in each prototype array—the motivation for this choice is discussed in Section 3.2. Due to the frequency dependence of the Galactic emission, as well as the antenna’s side lobes, it is expected that the absolute amplitude of the Galactic noise will be distinct for different frequency ranges. This is seen in the noise ranges of each plot in Figure 16, which are not the same but similar in magnitude. Nonetheless, we can state that both GP300 and G@A have detected the Galactic synchrotron emission. More detailed analyses are described in the references [44, 16].
@@ 7 Summary and perspectives
The Giant Radio Array for Neutrino Detection (GRAND) is a planned observatory of ultra-high-energy particles of cosmic origin: cosmic rays, neutrinos, and gamma rays. GRAND aims to detect the electromagnetic emission from air showers initiated by their interaction in the atmosphere using large ground arrays of antennas. The next large deployment goal of GRAND is to have two arrays consisting of 10,000 antennas each in the Northern and Southern Hemispheres. To achieve this, the GRAND Collaboration has adopted a staged approach to deployment, with three ongoing prototype arrays to test and validate the detection and reconstruction methods.
@@
Three prototype arrays were deployed in 2023. GRAND@Nançay, in France, is a 4-antenna array used as a testbench for the Collaboration’s European laboratories. GRANDProto300, in China, has gone through two stages of deployment, from an initial 13-antenna stage to its current 65-antenna stage. These stages act as pathfinders for the future final 300-antenna array. GRAND@Auger, in Argentina, stems from an agreement between the GRAND and the Pierre Auger Collaborations. The 10-antenna array was deployed on modified AERA stations. In this paper, we focus on the hardware, software, and early operation of GRANDProto300 and GRAND@Auger.
To adapt to local conditions and explore optimal designs for the final GRAND setup, the prototype array at each site featured a few different approaches to hardware and software. Throughout 2023, GRANDProto300 and GRAND@Auger were commissioned using daily data monitoring. The arrays have remained largely operational for about two years. Early data demonstrate that both systems can trigger on weak radio-frequency signals and are able to handle trigger rates up to 1 kHz. We provide a first characterization of the electromagnetic background and noise present at each site. The Galactic synchrotron emission is measured at both prototypes and will be used to refine the calibration of the GRAND detection units. In GP300, there are fewer RFI sources compared to G@A. Nonetheless, the configuration at each site is such that baseline levels of the PSDs are the same for both prototype arrays, around $10^{-2}$ mV2 MHz-1.
The GRAND Collaboration aims to achieve its goals while keeping its carbon footprint low [15]. Following the life cycle analysis of the current prototype GRAND DUs [53], we have identified two design modifications that will decrease the environmental impact of the experiment. First, changing the alloy composition of our mechanical structure’s stainless steel [53] would lower the impact by 66%. Secondly, using longer duration batteries would diminish the impact almost linearly. These design changes will be integrated into the preparation for the production of the next batch of antennas that will be deployed to complete the GRANDProto300 array.
Future stages of the GRAND experiment will face diverse environmental and electromagnetic background and noise conditions, due to the multiple geographical locations. The demonstrated adaptability of the GRAND hardware and software indicates the feasibility of combining data from the different arrays. This is essential for achieving full-sky coverage and large-scale exposure of future stages of GRAND. Presently, the data collected by GRANDProto300 and GRAND@Auger are being searched for air showers initiated by ultra-high-energy cosmic rays. In GRANDProto300, we have identified the first cosmic-ray candidates [29, 33, 2]. In GRAND@Auger, the detection of cosmic-ray candidates is validated through their coincident detection by the Pierre Auger Observatory and a first candidate has been observed [16].
@@ Acknowledgments
The GRAND Collaboration acknowledges the significant contributions to this manuscript of Beatriz de Errico, for her work on data management and analysis, site survey and construction, DAQ software debugging, manuscript writing, and figure preparation; and of Shen Wang, for his work on GP300 site survey and construction, hardware verification, and manuscript planning and writing. In addition, the Collaboration acknowledges significant contributions from the following Collaboration members: Mauricio Bustamante (coordination, manuscript writing, figures), Nicoleta Cucu Laurenciu (firmware development and trigger algorithm implementation), Bohao Duan (construction, GP300 DAQ software development), François Legrand (section on data storage), Pengxiong Ma (construction, RFI identification, GP300 data management and analysis), Frédéric Magnard (communications at GP300, DAQ and communications at Auger), Stavros Nonis (data analysis, RF chain implementation, and figures), Lech Piotrowski (GP300 DAQ software), Daniel Szálas Motesiczky (LNA design, electronics validation), Xishui Tian (laboratory trigger tests), Anne Timmermans (data monitoring web page), Charles Timmermans (site survey, construction, DAQ software development and debugging, data management and analysis, manuscript writing and revising, figures), Feng Wei (LNA design), Xin Xu (antenna response, RF chain optimization, data analysis), Xing Xu (construction, GP300 hardware verification, firmware development and trigger algorithm implementation), Pengfei Zhang (site survey, redesign and functional test of detection unit, construction, data analysis, RF chain generation), Yi Zhang (site survey and construction, data analysis, manuscript writing). The GRAND Collaboration is grateful to Arno Engels (Radboud University) for the mechanical design of the detection units, to Fei Gao (Xidian University) for that and for the thermal design of the detection units, to the local government of Dunhuang during site survey and deployment approval, to Yu Tang for his help on-site at the GRANDProto300 site, and to the Pierre Auger Collaboration, including the staff in Malargüe, for the warm welcome and continuing support, and the Auger Publications Committee, especially Silvia Mollerach, for their feedback. The GRAND Collaboration acknowledges the support from the following funding agencies and grants. Brazil: Conselho Nacional de Desenvolvimento Cientiífico e Tecnológico (CNPq); Fundação de Amparo à Pesquisa do Estado de Rio de Janeiro (FAPERJ); Coordenação Aperfeiçoamento de Pessoal de Nível Superior (CAPES). China: NAOC, National SKA Program of China (Grant No. 2020SKA0110200); National SKA Program of China (Grant No. 2025SKA0110100); Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061); Program for Innovative Talents and Entrepreneurs in Jiangsu, and High-end Foreign Expert Introduction Program in China (No. G2023061006L); China Scholarship Council (No. 202306010363). Denmark: Villum Fonden (project no. 29388). France: “Emergences” Programme of Sorbonne Université; France-China Particle Physics Laboratory; Programme National des Hautes Energies of INSU; for IAP—Agence Nationale de la Recherche (“APACHE” ANR-16-CE31-0001, ANR-23-CPJ1-0103-01), CNRS Programme IEA Argentine (“ASTRONU”, 303475), CNRS Programme Blanc MITI (“GRAND” 2023.1 268448), CNRS Programme AMORCE (“GRAND” 258540); Fulbright-France Programme; IAP+LPNHE—Programme National des Hautes Energies of CNRS/INSU with INP and IN2P3, co-funded by CEA and CNES; IAP+LPNHE+KIT—NuTRIG project, Agence Nationale de la Recherche (ANR-21-CE31-0025); IAP+VUB: PHC TOURNESOL programme 48705Z. Germany: NuTRIG project, Deutsche Forschungsgemeinschaft (DFG, Projektnummer 490843803); Helmholtz—OCPC Postdoc-Program. Poland: Polish National Agency for Academic Exchange within Polish Returns Program no. PPN/PPO/2020/1/00024/U/00001,174; National Science Centre Poland for NCN OPUS grant no. 2022/45/B/ST2/0288. USA: U.S. National Science Foundation under Grant No. 2418730. Computer simulations were performed using computing resources at the CC-IN2P3 Computing Centre (Lyon/Villeurbanne, France), partnership between CNRS/IN2P3 and CEA/DSM/Irfu, and computing resources supported by the Chinese Academy of Sciences.